peter invested 450 in insurance and stocks. He puts $50 more in insurance than than stocks. How much did he invest in each?
Answer by brysca(112)

So to put this in simplest terms you can say this, right?
x+y=450
x=stocks
y=insurance
In the question, they say that Peter invested $50 more in insurance than stocks so you can say
insurance=stocks+50
y=x+50
Now just substitute y in the original equation
x+y=450
x+(x+50)=450
2x+50=450
2x=400
x=200
Now substitute x in the insurance equation to get the value for insurance.
y=x+50
y=200+50
y=250
Therefore, Peter invested $200 into stocks and $250 into insurance.
-Bryan